a) The 82 kg Bogaraj left the ramp about 45 m lower than his starting position. (That is the difference in elevation was 45m) If you ignore the slowing forces of kinetic friction and wind drag,what was his kinetic energy at that point?

b) what was his speed at that point?

c) If Bogataj has simply fallen straight down 45 m rather than skiing down a ramp, what would be his kinetic energy?

d) After hitting the ground Bogataj crashed wildly and descended down the hill another 20 vertical meters before coming to rest. Here friction cannot be ignored! How much total work was done on Bogataj to bring him to rest?

e) In part a) you were told to ignore friction and air,but based on your knowledge of friction and drag,which one would be more important to account for?
Answer #1

a) Gravitational force is a conservative force and hence change in kinetic energy = work done by gravitational force

there is no friction force or air drag, therefore

mgh = 1/2 mv2

Kinetic energy = mgh = 82*9.8*45 = 36162 Joules

b)

mgh = 1/2 mv2

sqrt below will refer to square root of

v = sqrt(2gh) = sqrt(2*9.8*45) = 29.7 m/s

c) Same as a) 36162 Joules

d) Work done in bringing to rest = 36162+(82*9.8*20) = 52234 Joules

e) In a) friction would be more important as Bogaraj was skiing down the ramp therefore in contact with rough surface.